2fourth root of 3=fourth root of 48, how does that work out?
Answer by stanbon(75887)

2fourth root of 3=fourth root of 48, how does that work out?
---------------------------------------
That is true.
(48)^(1/4) = (16*3)^(1/4) = (16)^(1/4)*(3)^1/4) = 2*(3)^(1/4)
-------------------------------------------
Cheers,
Stan H.